### v1

# VidLaDA: Bidirectional Diffusion Large Language Models for Efficient Video Understanding

Abstract
Standard Autoregressive Video LLMs inevitably suffer from causal masking biases that hinder global spatiotemporal modeling, leading to suboptimal understanding efficiency. We propose VidLaDA, a Video LLM based on Diffusion Language Model utilizing bidirectional attention to capture bidirectional dependencies. To further tackle the inference bottleneck of diffusion decoding on massive video tokens, we introduce MARS-Cache. This framework accelerates inference by combining asynchronous visual cache refreshing with frame-wise chunk attention, effectively pruning redundancy while preserving global connectivity via anchor tokens. Extensive experiments show VidLaDA outperforms diffusion baselines and rivals state-of-the-art autoregressive models (e.g., Qwen2.5-VL and LLaVA-Video), with MARS-Cache delivering over 12× speedup without compromising reasoning accuracy. Code and checkpoints are open-sourced at https://github.com/ziHoHe/VidLaDA.
Machine Learning, ICML
Figure 1: The overall architecture of VidLaDA. Input video frames are encoded and spatially pooled (via $2\times 2$ downsampling) before being unrolled into a sequence of Spatiotemporal Visual Tokens $V$. These tokens, combined with the text prompt $P$ and the noised answer $X_{t}$, are processed by the Diffusion Language Model. Unlike autoregressive models, VidLaDA utilizes full bidirectional attention. This design enables global, unconstrained interactions both within and across visual and textual modalities, ultimately facilitating the parallel prediction of the target answer $X_{0}$.
1 Introduction
Video Large Language Models (Video LLMs) (Lin et al., 2024; Cheng et al., 2024) have recently emerged as a key paradigm for connecting visual perception with language reasoning, showing strong potential in tasks such as video captioning, spatiotemporal question answering, and long-horizon decision making. Most existing systems follow a standard recipe that couples a pretrained vision encoder (e.g., ViT (Zhu et al., 2023; Tschannen et al., 2025)) with an Autoregressive (AR) LLM (Grattafiori et al., 2024; Team et al., 2024). Video frames are first encoded into visual tokens and projected into the language embedding space, after which the model generates responses via next-token prediction under a causal attention mask (Lin et al., 2024).
While this AR-based paradigm has driven rapid progress, it potentially ignores the fundamental mismatch between AR and the spatiotemporal nature of video. Unlike text, visual semantics (objects, relations, and event cues) are distributed across space and time without an inherent left-to-right ordering (Yu & Wang, 2025). However, after rasterizing video tokens into a 1D sequence and feeding them into an AR decoder, the causal mask enforces a strictly unidirectional dependency. This creates an asymmetric receptive field where early tokens are visible to many subsequent positions while late tokens are structurally under-attended, inducing positional bias and non-uniform utilization of visual evidence. This results in suboptimal understanding efficiency, i.e., the Video LLM fails to fully exploit the available bidirectional visual information within its receptive field. We formalize this inefficiency issue from two complementary perspectives: (i) causal attention yields a visibility-frequency imbalance that encourages early receptive field behaviors (Proposition 3.1); and (ii) due to the restricted access to future visual tokens, the AR decoding stack admits a strictly lower upper bound on the usable information extracted from the bidirectional vision encoder (Proposition 3.2). Empirically, these biases manifest in two ways: brittle intra-frame performance under controlled relocation of high-information patches, and degraded temporal robustness in causal video question answering when key evidence is sparse or occurs at different positions in the video (see Figure 2).
Moreover, video understanding is increasingly shifting from short-horizon perception to complex spatiotemporal reasoning (Ye et al., 2025b; Liu et al., 2025; Feng et al., 2025). Compared to images, videos are both temporally redundant and dynamically evolving, which often requires models to efficiently aggregate and understand key spatiotemporal evidence across the video and form a long-form reasoning chain. However, the standard AR paradigm faces a dual efficiency bottleneck in this context. First, regarding understanding efficiency (modeling front), strictly unidirectional temporal modeling can hinder capturing global temporal relationships across different moments in a video (Guo et al., 2025b; Chen et al., 2024b), which limits the effective aggregation of spatiotemporal evidence. Second, regarding generation efficiency (inference), AR decoding is inherently serial, making latency scale linearly with generated tokens and limiting generation efficiency for long-form reasoning for complex spatiotemporal reasoning. Conversely, Diffusion Language Models (DLMs) offer a compelling alternative: generation is formulated as iterative denoising in a discrete space with full bidirectional attention (Nie et al., 2025), which removes the causal constraint and enables parallel prediction of multiple tokens per step (Wu et al., 2025). This bidirectional decoding is advantageous for video understanding, where key spatiotemporal evidence should aggregate visual context globally across all frames and patches rather than follow a forced causal path.
(a) Intra-Frame: Spatial Ordering.
(b) Inter-Frame: Temporal Context.
(c) Inter-Frame: Frame Sparsity.
Figure 2: Comparison of Spatiotemporal Robustness. (a) Performance vs. spatial location of high-norm tokens. DLM-Based VLM remains invariant, whereas AR degrades when salient features shift from the start. (b) Performance vs. temporal location of the key event. DLM-Based VLM demonstrates stability across the timeline, while AR baselines show significant volatility. (c) DLM-Based VLM maintains high accuracy with fewer frames, demonstrating superior aggregation of sparse evidence compared to AR models.
Despite these advantages, directly applying DLMs to long-form video introduces a new efficiency bottleneck. Unlike AR inference, where prefix key/value states can be cached, and only the newest token is processed, vanilla DLM inference recomputes bidirectional attention over the entire multimodal sequence at every decoding step. With thousands of video tokens, this results in prohibitive $\mathcal{O}(|V|^{2})$ attention cost per step, which can negate the speed benefits of parallel decoding. Moreover, generic DLM acceleration techniques developed for text-only settings (Wu et al., 2025) ignore multimodal heterogeneity: in our analysis, visual and textual states exhibit markedly different decoding dynamics, and inter-frame attention presents strong local temporal structure with a small set of globally influential anchor tokens. Treating all modalities and all layers identically, therefore wastes computation on repeatedly recomputing stable context, especially for video inputs.
To unlock bidirectional diffusion for efficient video understanding, we propose VidLaDA (Video-Language Diffusion with mAsking), a Video LLM built on the DLM backbone with bidirectional attention to maximize understanding efficiency, and MARS-Cache (Multi-modal Asynchronous Refreshing Strategy), an inference framework that prunes redundancy specific to multimodal diffusion to maximize generation efficiency. Trained via a multi-stage curriculum on a comprehensive video dataset incorporating our newly constructed collection, VidLaDA couples a strong vision encoder with DLM-based bidirectional decoding, enabling unconstrained global interactions across video tokens, prompts, and partially denoised answers. MARS-Cache accelerates long-video DLM inference by combining (i) frame-wise chunk attention to exploit temporal locality, reducing the visual refresh cost to linear $\mathcal{O}(|V|)$, (ii) adaptive anchor token searching to preserve necessary global connectivity, and (iii) modality- and depth-wise asynchronous cache refreshing to update stable visual context less frequently than dynamic textual states. Our main contributions are summarized as follows:
  1. 1.
    We introduce VidLaDA, the first family of DLM-Based Video LLMs. By utilizing bidirectional decoding, VidLaDA mitigates the asymmetric receptive field issue in causal attention and improves the theoretical upper bound of spatiotemporal understanding in Video LLMs.
  2. 2.
    VidLaDA consistently outperforms existing DLM-based baselines (e.g., LLaDA-V (You et al., 2025) and Dream-VL (Ye et al., 2025a)). Furthermore, it remains highly competitive with state-of-the-art open-sourced AR-based Video LLMs (e.g., Qwen2.5-VL (Bai et al., 2025b) and LLaVA-Video (Zhang et al., 2024b)).
  3. 3.
    We propose MARS-Cache, a Multi-modal Asynchronous Refreshing Strategy, motivated by observations of modality-wise stability and attention structure during decoding. Substantially, MARS-Cache achieves a more than $12\times$ throughput improvement compared to vanilla DLM without compromising reasoning accuracy.
2 Preliminary
Problem Formulation. Given a video input $\mathcal{V}$ and a textual prompt $P$, the goal of a Video Large Language Model is to generate a target response $R=\{r_{1},r_{2},\dots,r_{L}\}$. The video is typically processed into visual tokens $V$, which form the multimodal context alongside prompt textual tokens $P$.
From Autoregressive to Diffusion Modeling. Standard Video LLMs adopt an AR paradigm (Li et al., 2024a; Zhang et al., 2024b). They model the joint probability of the response as a product of conditional probabilities, optimized via the negative log-likelihood loss. Crucially, AR models rely on the causal attention, where the token can only attend to its preceding ones. This unidirectional dependency inherently prevents early tokens (visual or textual) from interacting with subsequent context, limiting the model’s ability to capture global spatiotemporal relationships.
In contrast, our VidLaDA is built upon Diffusion Language Models (DLMs) (Nie et al., 2025; Ye et al., 2025c), which treat generation as a bidirectional iterative denoising process. Instead of sequential prediction, DLMs utilize a Masked Diffusion Model (MDM) framework. During training, a subset of response tokens is randomly replaced by a special [MASK] token based on a timestep $t$. The model learns to predict the original identities of these masked tokens simultaneously (Nie et al., 2025). Moreover, the DLM architecture allows VidLaDA to employ a full bidirectional attention mechanism without causal masking. Consequently, the aggregation and prediction of any token depends on the entire global context (both unmasked and masked tokens in $V$, $P$, and $R$), facilitating unconstrained interaction between visual and textual modalities. For detailed mathematical formulations of the forward and reverse diffusion processes, please refer to Appendix A.
(a) Step 1/Layer 8
(b) Step 1/Layer 21
(c) Step 32/Layer 8
(d) Step 32/Layer 21
Figure 3: Visualization of Attention Patterns. We display the attention score matrices across different denoising steps and layers. The heatmaps reveal two distinct structural properties utilized by our MARS-Cache design: (1) Chunk-wise Locality, visible as diagonal blocks where tokens primarily attend to their temporal neighbors, and (2) Global Anchor Tokens, manifested as prominent vertical bands where specific tokens consistently attract global attention from the entire sequence, regardless of the diffusion step or network depth.
3 VidLaDA: Efficient Video Understanding
3.1 Why Bidirectional DLM for Video Understanding?
Proposition 3.1.
AR possesses an asymmetric receptive field that precludes uniform spatiotemporal processing.
Proof.
See Appendix C.1
Proposition 3.2.
The AR paradigm inherently limits the utilization of omnidirectional spatiotemporal features encoded by ViT, resulting in a strictly lower mutual information upper bound compared to a bidirectional decoding paradigm.
$\sum_{t=1}^{T}\sup I(z_{t}^{\text{bidi}};\mathcal{V})>\sum_{t=1}^{T}\sup I(z_{t}^{\text{uni}};\mathcal{V}),$ (1)
where $\mathcal{V}$ denotes the original visual inputs, and $z^{*}_{t}$ denotes the representation encoded by ViT and LLM at timestep $t$.
Proof.
See Appendix C.2
3.1.1 The Impact of Token Spatial Ordering
To empirically validate the asymmetry described in Proposition 3.1, we investigate how the spatial ordering of visual tokens affects semantic stability at the single-frame level.
In ViT, the patch order is conventionally rasterized (top-left to bottom-right). Inspired by (Luo et al., 2025), we use the $\ell_{2}$-norm of the ViT output features as a proxy for the information density of visual tokens. We identify the top-$k$ high-norm tokens and virtually relocate them within the sequence fed to the LLM, but maintaining the same positional IDs in sequence. This design isolates the effect of causal visibility in the decoder from changes in spatial positional encoding. We define a position ratio $r\in[0,1]$, where $r=0$ places high-norm tokens at the start of the visual sequence and $r=1$ pushes them to the end.
As illustrated in Figure 2(a), we evaluate the models on the MME (Yin et al., 2024) benchmark under this shuffling protocol (cf. Appendix F.2). The Bidirectional DLM-Based VLM (You et al., 2025) exhibits a near-constant performance profile (red line), demonstrating invariance to the causal position of semantic features.
Conversely, AR-based models (LLaVA-OneVision (Li et al., 2024a), LLaVA-1.5 (Liu et al., 2024b)) display performance degradation, particularly when high-information tokens are shifted towards the middle or end of the sequence. This corroborates Proposition 3.1: AR models over-rely on early tokens as a receptive field to maintain semantic interpretation. When salient features are displaced from the high-visibility start positions to later positions (where visibility is lower), the autoregressive attention mechanism fails to allocate sufficient attention mass to them, treating them as contextually less significant solely due to their temporal index.
3.1.2 The Impact of Event Position and Sparsity
To verify spatiotemporal robustness in realistic settings beyond artificial single-image shuffling, we evaluate on ReXTime (Chen et al., 2024a) (cf. Appendix F.3), where each question depends on a specific event segment. We bucket samples by the temporal position ratio of the ground-truth event and report accuracy under uniform frame sampling.
As shown in Figure 2(b), AR baselines exhibit a U-shaped sensitivity to event position. Since the original spatiotemporal structure is preserved (unlike the single-frame experiment), this trend reflects the interplay between causal masking and positional effects such as RoPE long-range attenuation (Su et al., 2024) and recency bias (Liu et al., 2024c). Specifically, events occurring early suffer from long-range attenuation under RoPE, while mid-sequence events ($40\%\!\sim\!80\%$) are particularly fragile: they are neither close enough to benefit from recency nor sufficiently early to act as the early receptive field under causal decoding (Proposition 3.1), making their evidence more likely to be diluted or overwritten during autoregressive abstraction.
Notably, LLaVA-Video (Zhang et al., 2024b) exhibits the most severe instability despite extensive video training compared to LLaVA-OneVision (Li et al., 2024a), suggesting that scaling data alone cannot remove the structural limitations imposed by the early receptive field of AR. This limitation becomes more evident when event evidence is sparse. As shown in Figure 2(c), AR models degrade more sharply as fewer event frames are available, indicating that answer-relevant information of the event is more likely to be dispersed during deep semantic abstraction once it is not continuously reinforced under causal attention. In contrast, the bidirectional DLM-based Video LLM maintains a near-flat accuracy profile across event positions (Figure 2(b)) and remains accurate even under sparse event frames (Figure 2(c)). This behavior is consistent with Proposition 3.2: bidirectional decoding allows answer-relevant representations to directly aggregate global spatiotemporal context across all frames. This mechanism enhances understanding efficiency and prevents the irreversible loss of mid-sequence event details during deep semantic abstraction, even under sparse evidence conditions. Furthermore, it inherently eliminates the visibility asymmetry that induces early receptive field in AR decoding (Proposition 3.1).
Figure 4: Modality-Dependent State Evolution. We visualize the cosine similarity matrix of hidden states across different inference steps. Left: Visual tokens exhibit high stability, indicating low drift during the decoding process. Right: Textual tokens show lower similarity between steps, indicating high volatility.
Figure 5: Depth-Dependent Hidden State Drift. The heatmap illustrates the magnitude of hidden state drift (measured as 1-Cosine Similarity) across network layers (Y-axis) and inference steps (X-axis). Shallow layers remain stable with minimal drift, whereas deep layers exhibit significant volatility.
3.2 Training Pipeline
Data Composition. To address the scarcity of training data for video understanding beyond 3 minutes, we first curate a dataset emphasizing 2-30 minute videos. The pipeline involves (1) temporal stratification, (2) instruction synthesis via LLM, (3) text-only bias filtering, and (4) VLM consistency voting (See Appendix D.1.1 for details).
Multi-Stage Training. We also employ a multi-stage training strategy for VidLaDA (See Appendix D.1.2 for details), including (1) short-clip temporal pre-training, (2) temporal scaling warm-up, and (3) long-form video expansion.
3.3 Model Architecture
The overall architecture of VidLaDA, as shown in Figure 1, integrates a robust vision encoder with a DLM backbone to enable holistic spatiotemporal-friendly reasoning.
Vision Encoder and Spatial Pooling. We utilize SigLip2-SO400M (Tschannen et al., 2025) to extract feature representations from the input video frames. To efficiently manage the extensive token sequence inherent to video data, we adopt a straightforward spatial pooling strategy common in recent Video LLMs (Zhang et al., 2024b; Li et al., 2024a). Specifically, after projection, the visual embeddings are reshaped to their original 2D spatial grid and downsampled via $2\times 2$ bilinear interpolation. This operation reduces the visual sequence length by a factor of 4, balancing computational efficiency with the preservation of spatial structure.
Bidirectional Diffusion Language Model. The core unit of VidLaDA is the Diffusion Language Model with the full bidirectional attention mechanism. This architecture addresses the limitations of traditional Autoregressive (AR) VLMs to some extent. Standard AR models rely on causal attention, which imposes a strict left-to-right dependency. This structure introduces the asymmetric receptive field, where early visual tokens are structurally prevented from attending to subsequent visual or textual context from user input, disrupting the global topology of the vision (See Section 3.1). In contrast, VidLaDA leverages a full bidirectional attention mechanism. This design eliminates the causal constraint, allowing every token, whether visual or textual, to attend to the holistic sequence simultaneously. This ensures that the vision encoder and language model are deeply coupled, preserving the integrity of dynamic video information.
Furthermore, unlike the serial, token-by-token generation of AR models, VidLaDA is capable of predicting multiple tokens simultaneously within a single step (Nie et al., 2025). This property significantly enhances throughput potential, making it better suited for the long-context reasoning requirements of complex spatiotemporal understanding tasks.
4 Efficient Video Reasoning via MARS-Cache
While DLMs offer the advantage of parallel decoding, they fundamentally differ from AR models during inference. In AR generation, past key-value pairs are cached, restricting computation to the newest token. In contrast, the bidirectional nature of vanilla DLMs necessitates re-computing attention for the entire sequence at every decoding step (Nie et al., 2025; Ye et al., 2025c; You et al., 2025). Consequently, for video understanding, the massive number of video tokens must be processed iteratively alongside the text, creating a prohibitive computational burden, thereby partially offsetting the efficiency advantages of parallel decoding. To address this, we propose the Multi-modal Asynchronous Refreshing Strategy (MARS-Cache), a framework designed to prune redundancy based on the spatiotemporal behavior of DLM-Based Video LLMs.
4.1 Empirical Observations
We analyze the internal state evolution during the decoding process and identify four distinct observations: (1) Chunk-wise Locality with Global Anchor Tokens. As shown in Figure 3, inter-frame interactions primarily exhibit local dependencies. However, a specific subset of anchor tokens consistently attracts high attention scores globally. Crucially, we observe that the spatial positions of these anchor tokens remain stable across diffusion steps, indicating that the global information hubs are determined early in the inference process. Furthermore, we observe a hierarchical inclusion property: the set of anchor tokens in deeper layers is a subset of those in shallow layers ($\mathcal{I}_{\text{deep}}\subset\mathcal{I}_{\text{shallow}}$), adhering to a dense-to-sparse pyramid structure. (2) Modality-Dependent Drift. As shown in Figure 5, there is a notable disparity in hidden state evolution between modalities. Text hidden states exhibit higher temporal drift across steps compared to vision tokens. (3) Depth-Dependent Stability Variance. As shown in Figure 5, hidden state drift is not uniform across network depth. Shallow layers exhibit high stability, while deep layers show significant volatility. (4) Progressive Attention Sparsity. As shown in Figure 6, the attention distribution evolves through the network depth, transitioning from a near-uniform distribution in shallow layers to a highly peaked, sparse distribution in deep layers.
Figure 6: Visualization of Progressive Attention Sparsity. We display the aggregated attention maps of visual tokens from the entire sequence at varying network depths. The distribution transitions from a diffuse, global pattern in shallow layers to a highly peaked, semantically focused pattern in deep layers.
Table 1: Comparison with SOTA Video LLMs across comprehensive benchmarks. The models are categorized into AR-Based and DLM-Based models. #S: Model Size (Parameters). #F: Number of input frames. Models marked with * denote the reproduced results.
Model #S #F Video-MMMU LongVideoBench LVBench EgoSchema MVBench $_{\text{dev}} MLVU$$_{\text{test}}$ Video-MME
AR-Based Video LLM
VideoLLaMA2.1 (Cheng et al., 2024) 7B 32 - - 36.2 53.1 57.3 61.2 - 54.9
VideoChat2 (Li et al., 2025a) 7B 16 - 36.0 - 54.4 - - - 39.5
InternVL2.5 (Chen et al., 2024d) 7B 64 - 60.0 38.4 51.5 72.0 68.9 - 64.2
Qwen2-VL (Wang et al., 2024a) 7B 2fps - - - 66.7 67.0 - - 63.3
Qwen2.5-VL* (Bai et al., 2025b) 7B 64 47.4 - 45.3 65.0 69.6 62.8 45.3 63.9
Video-LLaVA (Lin et al., 2024) 7B 8 - - - 38.4 - 47.3 - 39.9
LLaVA-NeXT-Video (Zhang et al., 2024a) 7B 32 - 43.5 - 43.9 33.7 - - 46.5
LLaVA-OneVision* (Li et al., 2024a) 7B 32 33.9 56.5 - 60.1 56.7 64.7 45.3 58.5
LLaVA-Video* (Zhang et al., 2024b) 7B 64 37.1 58.2 41.5 57.3 58.6 70.8 50.4 63.7
DLM-Based Video LLM
Dream-VL (Ye et al., 2025a) 7B - - - - - - 61.1 - 61.5
SDAR-VL (Cheng et al., 2025) 8B - - - - - - 65.0 - 60.8
LLaDA-V* (You et al., 2025) 8B 32 43.3 58.6 36.4 57.9 53.1 59.4 44.1 56.4
VidLaDA (Ours) 8B 64 46.6 61.4 44.7 64.5 59.4 69.2 53.4 64.2
4.2 Multi-modal Asynchronous Refreshing Strategy
Multi-modal Asynchronous Refreshing. Leveraging the drift disparities (Obs. 2 & 3), we introduce a hierarchical asynchronous refresh schedule. This strategy applies to context hidden states (visual tokens, prompts, decoded text).
We maintain caches for context hidden states and skip updates based on the refresh interval. We partition model layers into groups (for simplicity, we use 4 groups in this work) and assign a grouped refresh interval $\tau_{(g,m)}$: (i) Modality-wise: Since text drifts more (Obs. 2), we set $\tau_{(*,t)}<\tau_{(*,v)}$, refreshing visual caches less frequently. (ii) Layer-wise: Deep layers are more volatile (Obs. 3), so we assign smaller intervals to deeper groups (i.e., $\tau_{(g-1,*)}>\tau_{(g,*)}$).
Formally, a hidden state $H^{t,g}$ is updated only if $t\mod\tau_{(g,m)}=0$; otherwise, the cached state is reused. To preserve feature consistency across depths, we constrain the refresh interval of a shallow group to be a fixed integer multiple of the deeper group (i.e., $\tau_{(g-1,*)}=k\cdot\tau_{(g,*)}$). This enforces a synchronization constraint: whenever a stable shallow layer is updated, the volatile deep layers are simultaneously refreshed, ensuring that low-level features are always synchronized with high-level abstractions.
Frame-wise Chunk Attention. While asynchronous refresh reduces the number of updates of visual tokens during inference, the update overhead of full attention for massive visual tokens still creates a major computational bottleneck when the visual cache requires refreshing. Based on Obs. 1, we observe that video tokens primarily attend to their temporal neighbors. Motivated by this sparsity, we adopt Frame-wise Chunk Attention to accelerate the inference.
Formally, we define $F_{n}$ as the subset of visual tokens belonging to the $n$-th frame. Accordingly, the local temporal neighborhood for frame $n$ is defined as $\mathcal{N}(F_{n})=F_{n-1}\cup F_{n}\cup F_{n+1}$. The attention for visual tokens within frame $n$ is then computed as:
$O_{[F_{n}]}=\text{Softmax}\left(\frac{Q_{[F_{n}]}K_{[\mathcal{N}(F_{n}),P]}^{\top}}{\sqrt{d_{k}}}\right)V_{[\mathcal{N}(F_{n}),P]},$ (2)
where $P$ denote the sets of text tokens of the prompt, $d_{k}$ represents the dimensionality of the key, $O_{[F_{n}]}$ represents the attention outputs for frame $n$, and the subscript $\mathcal{N}(F_{n}),P]$ denotes the concatenation of tokens from the neighboring frames and the text sequence. This operation effectively reduces the total complexity of the visual attention component from quadratic $\mathcal{O}(|V|^{2})$ to linear $\mathcal{O}(|\mathcal{N}(F_{n})|\times|V|)$.
Moreover, to ensure the robustness of the visual representation, we still perform full attention at the first decoding step to establish precise global context and initialize the visual cache. In subsequent steps, when the visual cache requires refreshing, we switch to the efficient chunk attention.
Adaptive Anchor Token Searching. However, naively restricting global interactions through chunking compromises the long-range information flow, leading to severe performance degradation (as analyzed in Appendix E.4.1). We hypothesize that specific anchor tokens within each frame act as critical inter-frame bridges, aggregating and transmitting global visual semantics across the temporal sequence. Exclusively relying on frame-wise chunk attention impedes these pathways, disrupting the hierarchical propagation of visual information, particularly in deeper layers. To restore global connectivity efficiently, we propose an optimized searching strategy comprising the following key components: (i) Adaptive Proxy Scoring: Computing the full attention map to identify anchor tokens is cost-prohibitive. Instead, we employ equidistant subsampling. Specifically, we sample a small subset of query tokens $\hat{S}$ (e.g., 32) to compute a low-rank proxy attention matrix:
$\hat{A}=\text{Softmax}(Q_{[\hat{S}]}K_{[V]}^{\top}/\sqrt{d_{k}})\in\mathbb{R}^{|\hat{S}|\times|V|}.$ (3)
To mitigate self-attention biases, we mask the diagonal entries where queries attend to themselves, yielding the debiased attention matrix $A$. (ii) Temporal Reuse: Leveraging Obs. 1, we perform this proxy scoring only at the first decoding step ($t=1$). The identified sink indices are cached and reused for all subsequent sparse attention steps, reducing the extra computational overhead. (iii) Group-wise Allocation: Leveraging the hierarchical consistency of anchor tokens in Obs. 1, we avoid per-layer searching. Instead, we search once per layer group. We allocate a decreasing budget $k_{g}$ for deeper groups (e.g., $k_{g-1}>k_{g}>\dots$). For group $g$, we aggregate the importance scores from the proxy matrix $A_{g_{\text{start}}}$ (computed at the group’s first layer) and select the top-$k_{g}$ indices.
To ensure a contiguous memory layout and consistent computational load, we enforce a fixed budget of anchor tokens per frame. For each frame $F_{j}$, we aggregate the importance scores across the sampled queries. Consequently, at step $t=1$, the anchor indices $\mathcal{I}_{(g,F_{j})}$ for each frame within group $g$ are determined by:
$A^{\prime}=\sum_{i}A_{g_{\text{start}}}[:,F_{j}]\in\mathbb{R}^{|F_{j}|},\,\mathcal{I}_{(g,F_{j})}=\text{Top}_{k_{g}}\left(A^{\prime}\right).$ (4)
The final set of global anchor tokens, $\mathcal{I}_{g}=\bigcup_{j}\mathcal{I}_{(g,F_{j})}$, is subsequently made visible to all tokens in the sequence, while these anchor tokens retain the ability to attend to all other tokens, effectively restoring global information propagation. Finally, exploiting the permutation invariance of bidirectional attention, we can relocate the anchor tokens within each frame to the beginning of the sequence to optimize memory access patterns.
5 Experiments
5.1 Experimental Setup
In this section, we first evaluate VidLaDA against state-of-the-art (SOTA) AR- and DLM-Based Video LLMs across diverse benchmarks. Then, we investigate the performance of CoT inference combined with MARS-Cache. Finally, this section concludes with a series of ablation studies. Detailed experimental setups are provided in Appendix F.1.
5.2 Main Results
We benchmark VidLaDA against current leading open-sourced AR-Based models and SOTA DLM baselines. The comparative results are presented in Table 1.
As the first DLM-based Video LLM optimized with bidirectional spatiotemporal attention, VidLaDA establishes a new baseline for non-autoregressive video understanding. It consistently outperforms existing DLM approaches across all evaluated dimensions (LLaDA-V (You et al., 2025), SDAR-VL (Cheng et al., 2025), Dream-VL (Ye et al., 2025a)).
Furthermore, VidLaDA demonstrates remarkable competitiveness against top-tier open-sourced AR-based Video LLMs, including Qwen2.5-VL (Bai et al., 2025b), LLaVA-OneVision (Li et al., 2024a), and LLaVA-Video (Zhang et al., 2024b). Notably, VidLaDA exhibits strengths particularly in tasks necessitating complex spatiotemporal understanding, where unidirectional modeling may constrain the aggregation of distributed visual evidence. For instance, it surpasses LLaVA-Video and LLaVA-OneVision on LongVideoBench and outperforms the robust Qwen2.5-VL on MLVU (both Dev and Test splits). These results support the hypothesis that our bidirectional attention mechanism effectively addresses the asymmetric receptive field limitations of AR architectures by enabling more robust global dependency modeling (See Section 3.1).
Table 2: Evaluation of MARS-Cache Acceleration on Video Reasoning. We assess the generalizability and efficiency of MARS-Cache by applying it under Chain-of-Thought (CoT) settings. TPS denotes Throughput (Tokens Per Second).
[…]
Acc TPS Acc TPS Acc TPS
LLaVA-OV (CoT) 62.8 27.0 43.3 27.2 57.5 25.0
w/o KV-Cache - 0.7 - 0.7 - 0.8
LLaDA-V (CoT) 63.4 3.5 42.5 4.3 55.7 4.5
+ Dual-Cache 64.2 23.8 43.2 28.3 57.1 27.2
+ MARS-Cache 63.4 35.2 43.6 40.0 57.7 37.9
VidLaDA (CoT) 67.4 2.7 49.0 3.3 59.8 2.1
+ Dual-Cache 67.2 23.7 50.5 26.0 59.4 18.5
+ MARS-Cache 67.0 33.6 50.7 33.6 59.7 25.2
5.3 Results under CoT Inference
To unlock the complex reasoning capabilities of models, we employ a structured Chain-of-Thought (CoT) inference pipeline inspired by (Zhou et al., 2024). This pipeline proceeds through four distinct stages: (1) task prompt routing, (2) reasoning analysis, (3) self-reflection, and (4) final answer generation. Further details on the specific prompt designs and experimental setup are provided in Appendix D.2 and Appendix F.4, respectively.
We evaluate the efficiency and robustness of our proposed strategies on EgoSchema (subset), MLVU (test), and LongVideoBench. Quantitative results are summarized in Table 2, with the batch size set to 1. To ensure practical inference speeds and establish a strong baseline, we employ parallel decoding (Wu et al., 2025) by default across all DLM settings. The primary objective of this experiment is to verify the generalizability of MARS-Cache on DLMs (LLaDA-V and VidLaDA) under CoT inference, and to benchmark its efficiency against the AR baseline.
[…]
Overall, these results demonstrate that MARS-Cache substantially accelerates CoT-style video reasoning, generalizes well across DLM-Based Video LLMs, and achieves competitive (or even better) efficiency than the strong AR baseline while maintaining comparable reasoning performance.
5.4 Ablation Studies
In this section, we summarize the key conclusions from our ablation studies on the MARS-Cache design. Comprehensive experimental results are provided in Appendix E.4. (i) The Necessity of Anchor Token: Our assessment of global connectivity reveals that retaining anchor tokens is crucial; completely reverting to frame-wise chunk attention significantly degrades performance. We find that preserving full attention in shallow layers to maintain structural integrity while aggressively pruning anchors in deep layers achieves the optimal trade-off between reasoning accuracy and computational cost. (ii) Asynchronous Refreshing Strategy: Leveraging the observation that visual hidden states exhibit higher temporal stability than textual ones, we implement a modality-wise refreshing strategy. We conclude that setting the visual refresh interval larger than the textual interval (optimally $R_{v/t}\approx 2$) significantly boosts throughput without compromising accuracy, whereas excessive ratios lead to degradation. Additionally, regarding network depth, a pyramid-style schedule (increasing update frequency for deeper, more volatile layers) consistently outperforms uniform schedules and is orthogonal to the modality-wise refreshing strategy. (iii) Search Overhead: For the adaptive anchor searching mechanism, we determine that a query subset size of 128 offers a robust compromise. It is sufficient to identify high-quality global information hubs while keeping the computational overhead negligible, whereas using the full sequence leads to memory bottlenecks.
6 Conclusion
In this work, we introduce VidLaDA, a bidirectional diffusion language model framework that overcomes the spatiotemporal limitations of AR-Based Video LLMs. By enabling global bidirectional attention, VidLaDA overcomes the inherent limitations of causal decoding, allowing more efficient and balanced aggregation of spatiotemporal visual evidence for robust video understanding. To mitigate the computational overhead of DLM-Based Video LLM decoding, we further introduce MARS-Cache, an asynchronous cache refreshing strategy that exploits the stability of visual features and frame-wise locality to substantially accelerate inference. Extensive experiments demonstrate that VidLaDA outperforms existing DLM baselines and competes with SOTA AR-Based models, establishing a new, efficient paradigm for non-autoregressive video understanding.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Detailed Problem Formulation and Background
In this section, we provide the mathematical formulation for the Autoregressive baseline and the Diffusion Language Model backbone used in VidLaDA.
A.1 Autoregressive-Based Video Language Models
The prevailing paradigm for MLLMs treats video understanding as a conditional sequence generation task. Given visual tokens $V$ (derived from video $\mathcal{V}$ via a ViT and projector) and prompt tokens $P$, the model maximizes the conditional likelihood of the response $R$ autoregressively:
$\mathcal{L}_{\text{AR}}=-\sum_{i=1}^{|R|}\log p_{\theta}(r_{i}\mid V,P,R_{<i}),$ (5)
where $p_{\theta}$ is parameterized by a Transformer Decoder (Vaswani et al., 2017; Grattafiori et al., 2024). This approach necessitates a causal attention mask $M$, defined as:
$M_{ij}=\begin{cases}0&\text{if }j\leq i,\\ -\infty&\text{otherwise.}\end{cases}$ (6)
This factorization imposes a strictly unidirectional dependency, preventing visual or early text tokens from attending to future tokens.
A.2 Diffusion Language Models (DLMs)
DLMs (Nie et al., 2025; Ye et al., 2025c) formulate generation as a discrete diffusion process.
Forward Process. The forward process adds noise by independently masking tokens. Let $x_{0}$ denote the clean sequence (containing conditions $V,P$, and response $R$). At continuous time $t\in[0,1]$, each token $x_{t}^{(i)}$ in the response is replaced by [MASK] with probability $t$. The transition distribution is:
$q_{t|0}(x_{t}\mid x_{0})=\prod_{i}q_{t|0}(x_{t}^{(i)}\mid x_{0}^{(i)}),$ (7)
$\text{where }q_{t|0}(x_{t}^{(i)}\mid x_{0}^{(i)})=\begin{cases}1-t&\text{if }x_{t}^{(i)}=x_{0}^{(i)},\\ t&\text{if }x_{t}^{(i)}=\texttt{[MASK]}.\end{cases}$ (8)
Reverse Process and Objective. The reverse process aims to recover $x_{0}$ from the masked state $x_{t}$. A neural network $p_{\theta}(\cdot\mid x_{t})$ predicts the original tokens for masked positions. In visual instruction tuning, we condition on unmasked $V$ and $P$, applying masking only to $R$. The objective minimizes the variational upper bound:
$\mathcal{L}_{\text{DLM}}=-\mathbb{E}_{t,R_{0},R_{t}}\left[\frac{1}{t}\sum_{i=1}^{|R|}m_{t}^{(i)}\log p_{\theta}(R_{0}^{(i)}\mid V,P,R_{t})\right],$ (9)
where $m_{t}^{(i)}=\mathbb{I}[R_{t}^{(i)}=\texttt{[MASK]}]$ is the mask indicator.
Bidirectional Attention. Unlike AR models, the prediction of $R_{0}^{(i)}$ and understanding of $V,P,R$ in DLMs depend on the global context. Thus, we utilize full bidirectional attention:
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V.$ (10)
This mechanism enables global spatiotemporal modeling for $V$. During inference, the model starts from a fully masked $R_{1}$ and iteratively refines the sequence by predicting $R_{0}$ and re-masking based on a schedule.
Appendix B Related Work
B.1 Autoregressive-Based Video Large Language Models
Following the success of Large Language Models (LLMs) in text generation, Multimodal Large Language Models (MLLMs) have rapidly evolved to perceive visual inputs. Early works like LLaVA (Liu et al., 2024b) projected static visual features into the LLM’s token space, enabling image-text question answering. This paradigm was quickly adapted to the video domain, deriving the Video Large Language Models (Video LLMs). Models such as Video-LLaVA (Lin et al., 2024), VideoChat (Li et al., 2025a), VideoLLaMA (Cheng et al., 2024), LongVILA (Chen et al., 2024c), InternVideo (Wang et al., 2024b), mPLUG-Owl (Ye et al., 2024), and PLLaVA (Xu et al., 2024) aggregate temporal frames using pooling or Q-Former (Li et al., 2023a) structures before feeding them into an autoregressive LLM decoder. More recent state-of-the-art models, including LLaVA-OneVision (Li et al., 2024a), LLaVA-Video (Zhang et al., 2024b), Qwen2-VL (Wang et al., 2024a), and Qwen2.5-VL (Bai et al., 2025b) scale up the visual resolution and context length, achieving impressive performance on general video benchmarks.
However, these models fundamentally rely on the Autoregressive (AR) Model, based on the causal masking mechanism. As analyzed in Section 3.1, causal masking imposes a unidirectional dependency where visual tokens cannot attend to subsequent context during the encoding of the history. This structural asymmetry limits the modeling of global spatiotemporal dependencies and often leads to context fading in video understanding. VidLaDA addresses this by replacing the AR backbone with a bidirectional diffusion model, ensuring omnidirectional information flow.
B.2 Diffusion and Non-Autoregressive Language Models
Diffusion models have achieved remarkable success in continuous data generation (Peebles & Xie, 2023; Polyak et al., 2024; Li et al., 2023b). Recently, substantial progress has been made in adapting diffusion to discrete language modeling, including diffusion in continuous embedding space (Li et al., 2022) and masked/discrete formulations such as D3PM (Ho et al., 2020) and SSD-LM (Han et al., 2023). More recent large-scale Diffusion Language Models (DLMs), e.g., LLaDA (Nie et al., 2025), LLaDA-MoE (Zhu et al., 2025), LLaDA2 (Bie et al., 2025), and Dream (Ye et al., 2025c), demonstrate competitive generation quality to strong autoregressive LLMs while enabling parallel decoding via iterative denoising.
In multimodal settings, several works extend DLMs to vision-language tasks by conditioning on spatiotemporal tokens and largely retaining the standard MLLM pipeline (vision encoder $\rightarrow$ projector $\rightarrow$ language backbone), where diffusion is introduced mainly by swapping the AR backbone with a masked-denoising Transformer (e.g., LLaDA-V (You et al., 2025), Dream-VL (Ye et al., 2025a), and concurrent efforts (Li et al., 2025b; Cheng et al., 2025; Xin et al., 2025)). While these studies validate the feasibility of diffusion-based VLMs, they typically provide limited task- and modality-specific motivation or analysis for “Why DLM is more suitable than AR-Based model in image/video understanding?”, and they do not explicitly target the spatiotemporal challenges unique to video understanding.
VidLaDA bridges diffusion modeling and video understanding in a more video-oriented manner. We first analyze and justify why bidirectional DLM is suitable for spatiotemporal understanding in videos, beyond treating diffusion as a drop-in replacement for AR decoding. Building on this motivation, we present (to the best of our knowledge) the first DLM-based Video LLM that demonstrates strong performance on complex video understanding benchmarks. Finally, to make diffusion inference practical under long-form video contexts with massive video tokens, we introduce MARS-Cache, a multimodal, structure-aware caching framework that significantly reduces redundant computation during decoding.
Appendix C Proof
C.1 Proof of Proposition 3.1
Consider a sequence $X=\{x_{1},\dots,x_{T}\}$ processed by a causal attention mechanism. The attention score $A_{i,j}$ (the attention $x_{i}$ pays to $x_{j}$) is valid only if $j\leq i$. This constraint is enforced by a lower-triangular causal mask $M\in\{0,-\infty\}^{T\times T}$.
We define the visibility frequency $\mathcal{C}(x_{j})$ of a token $x_{j}$ as the number of times it acts as a key/value for other tokens in the sequence during a forward pass:
$\mathcal{C}(x_{j})=\sum_{t=1}^{T}\mathbb{I}(j\leq t)=T-j+1,$ (11)
where $\mathbb{I}$ is the indicator function.
For the initial token $x_{1}$, $\mathcal{C}(x_{1})=T$, meaning it is visible to and attended by every token in the sequence. In contrast, for a later token $x_{k}$ (where $k\gg 1$), $\mathcal{C}(x_{k})$ is significantly smaller.
Because the softmax function $\sigma(x_{j})=e^{x_{j}}/\sum_{k}e^{x_{k}}$ forces weights to sum to 1, the model requires a stable token to absorb excess attention mass when current semantic features are ambiguous (Xiao et al., 2024). Since $x_{1}$ is the only universally accessible token with maximum visibility frequency, the model learns an inductive bias to utilize $x_{1}$ as a specialized asymmetric receptive field.
For the visual sequence, since the index $j=1$ corresponds to a specific spatial location (e.g., top-left corner) due to rasterization, the learned bias induces a spatially non-uniform importance distribution: the visual start token is disproportionately weighted solely due to its position in the input sequence, thereby shaping the early receptive field.
C.2 Proof of Proposition 3.2
We model the information flow within the VLM as a Markov chain with the $M$-layer ViT and the $N$-layer LLM. Let the ViT map the original visual inputs $\mathcal{V}$ to an intermediate sequence of features $H^{(M)}=\{h_{1},\dots,h_{T}\}$. Since the encoder is bidirectional, $H^{(M)}$ captures global visual context. However, the subsequent processing in the $N$-layer LLM introduces a divergence.
Consider the generation of the latent representation $z_{t}$ at step $t\in\{1,\dots,T\}$ in the final layer.
According to the Data Processing Inequality, for any Markov chain $X\to Y\to Z$, we have $I(X;Z)\leq I(X;Y)$.
For bidirectional model $\mathcal{M}_{\text{bidi}}$, the generation of $z_{t}^{\text{bidi}}$ can access the entire intermediate feature set $H^{(M)}$ via full attention. The dependency is $\mathcal{V}\to H^{(M)}\to z_{t}^{\text{bidi}}$. Thus, the information capacity is bounded by the total encoded information:
$I(z_{t}^{\text{bidi}};\mathcal{V})\leq I(H^{(M)};\mathcal{V})$ (12)
For AR model $\mathcal{M}_{\text{uni}}$, due to the unidirectional causal mask, the generation of $z_{t}^{\text{uni}}$ is restricted to the current and past features $H_{\leq t}^{(M)}=\{h_{1},\dots,h_{t}\}$. The dependency is $\mathcal{V}\to H_{\leq t}^{(M)}\to z_{t}^{\text{uni}}$. Thus, the bound is:
$I(z_{t}^{\text{uni}};\mathcal{V})\leq I(H_{\leq t}^{(M)};\mathcal{V})$ (13)
Visual information is spatially distributed. The intermediate features $H^{(M)}$ are lossy compressions of $\mathcal{V}$ with spatial specialization. Consequently, the future tokens $H_{>t}^{(M)}$ contain independent information about $\mathcal{V}$ that is not fully captured by $H_{\leq t}^{(M)}$. This implies a strictly positive conditional mutual information:
$\small\Delta I_{t}=I(H^{(M)};\mathcal{V})-I(H_{\leq t}^{(M)};\mathcal{V})=I(H_{>t}^{(M)};\mathcal{V}\mid H_{\leq t}^{(M)})>0$ (14)
Therefore, for any $t<T$, the latent representation in the AR model has a strictly lower information upper bound than that of the bidirectional model.
We simply assume the final response $R$ is generated based on the probability $P(R\mid Z)$. The quality of $R$ in VLM is constrained by the joint mutual information of the entire sequence $I(R;\mathcal{V})$. Summing the pointwise information capacities (assuming independence for the upper bound estimation):
$\sum_{t=1}^{T}\sup I(z_{t}^{\text{bidi}};\mathcal{V})-\sum_{t=1}^{T}\sup I(z_{t}^{\text{uni}};\mathcal{V})=\sum_{t=1}^{T-1}\Delta I_{t}>0$ (15)
This confirms that the latent sequence $Z^{\text{uni}}$ essentially discards the future visual context during the deep semantic abstraction process (layers $M+1$ to $M+N$), leading to an irreversible loss of visual understanding capability compared to $Z^{\text{bidi}}$.
Appendix D Method Details
D.1 Training Pipeline
D.1.1 Data Composition
To remedy the lack of minute-scale, long-horizon temporal reasoning in existing video understanding training data, we construct a corresponding dataset through a robust construction pipeline. Utilizing FineVideo (Farré et al., 2024) as the seed source, we curate a high-quality dataset specifically emphasizing videos ranging from 2 to 30 minutes. Our pipeline leverages Deepseek-V3.1 (Liu et al., 2024a) as the LLM for instruction synthesis and text-based filtering, and Qwen3-VL-235B-A22B (Bai et al., 2025a) as the VLM for visual consistency verification. The pipeline consists of four distinct stages:
Data Acquisition and Temporal Stratification. Initial curation involves rigorous filtering for corruption and availability. To ensure balanced temporal coverage, we stratify the videos into five duration buckets (0-30s, 30s-60s, 1m-2m, 2-10m, 10-30m), with a strategic focus on processing the underrepresented 2-30 minute long-form intervals.
Automated Instruction Synthesis. Leveraging the rich metadata associated with the videos from FineVideo, we employ Deepseek-V3.1 to synthesize a diverse set of instruction-following tasks with the prompts from LLaVA-Video (Zhang et al., 2024b). These include Multiple Choice Questions (MCQ) and Open-ended QA, aiming to capture both fine-grained visual details and high-level narrative comprehension.
De-biasing via Text-Only Filtering. To mitigate text-only bias, where questions can be answered solely via subtitles or audio transcripts (e.g., news anchors, podcasts), we implement a filtering mechanism that only uses LLM to process data. We feed the generated questions into the LLM without visual inputs. Samples where the LLM correctly predicts the answer are discarded. This step effectively removes low-visual-information samples and ensures that the dataset strictly requires visual perception for reasoning.
Quality Assurance via Consistency Voting. To guarantee label reliability and minimize hallucinations, we employ a self-consistency voting strategy. The VLM generates responses for each sample three times under varying temperatures. An LLM evaluator then compares these responses against the ground truth. Only samples where the VLM achieves a consistency consensus (i.e., at least 2 out of 3 responses match the ground truth) are retained for the final training set.
Beyond the newly curated long-form video data, we integrate established high-quality benchmarks to ensure robust performance across diverse modalities. Specifically, we adopt LLaVA-Video-178K (Zhang et al., 2024b) as the foundational source for short-to-medium video instruction tuning, ensuring coverage of rich temporal dynamics and general video understanding. Furthermore, to maintain and enhance fine-grained spatial reasoning capabilities, we incorporate the MAmmoTH-VL (Guo et al., 2025a) dataset for image-based instruction following. This holistic data composition, which spans static images, short clips, and our proposed long-form videos, forms the basis of our unified training curriculum.
D.1.2 Multi-Stage Training Strategy
Table 3: Detailed Specifications of the Multi-Stage Training Curriculum. We outline the data composition, temporal resolution scaling, and optimization configurations across three stages, transitioning from short-clip alignment to long-form video reasoning.
Stage 1 Stage 2 Stage 3
#Frames 32 64 64
Max Sequence Length 8K 16K 16K
Video Sources Finevideo (Farré et al., 2024), ActivityNet-QA (Yu et al., 2019), NextQA (Xiao et al., 2021), Youtube (Zhu et al., 2023), ShareGPTVideo (Zhang et al., 2025), Ego4D (Grauman et al., 2022), PerceptionTest (Patraucean et al., 2023), Charades (Sigurdsson et al., 2016), YouCook2 (Zhou et al., 2018)
#Sample 1.8M 500K 500K
Min Duration $\sim$10s $\sim$1min $\sim$1min
Max Duration $\sim$3min $\sim$3min >>30min
Image Ratio 10% 10% 10%
Text Ratio 0% 0% 10%
Trainable Modules ViT/MLP/LLM ViT/MLP/LLM MLP/LLM
Learning Rate 2e–6/1e–5/1e–5 2e–6/1e–5/1e–5 2e–6/2e–6
We employ a three-stage curriculum learning strategy with the initial checkpoint from LLaDA-V (You et al., 2025), focusing on effectively adapting the model from static image to long-form video understanding. This approach progressively scales the temporal resolution and context length, ensuring stable convergence while mitigating the catastrophic forgetting of short-term temporal dynamics. The overall recipe is listed in the Table 3.
Stage 1: Short-Clip Temporal Pre-Training. The primary objective of the initial stage is to equip the static VLM with fundamental temporal perception capabilities. We utilize a large-scale collection of 1.8M short videos and images, with durations spanning from approximately 10 seconds to 3 minutes, sourced from diverse datasets. In this phase, the model is trained with a temporal resolution of 32 frames and a maximum sequence length of 8K tokens. We perform full-parameter fine-tuning on the Vision Transformer (ViT), Projector (MLP), and LLM backbone to align the vision encoder’s temporal features with the language space using learning rates of 2e-6, 1e-5, and 1e-5, respectively.
Stage 2: Temporal Scaling Warm-up. To bridge the gap between the short and long durations of the same video, we utilize an intermediate warm-up stage that transitions the model to higher temporal resolutions. We select 500K samples from the same dataset in Stage 1, and double the temporal sampling density to 64 frames while expanding the context window to 16K tokens. We continue with full-parameter tuning using the same learning rates as Stage 1, but shift the focus to stabilizing the adaptation to longer sequence dependencies. This stage is used to prevent performance degradation on short-term actions while preparing the attention mechanism for extended temporal spans.
Stage 3: Long-Form Video Expansion. The final stage focuses on reasoning over long and ultra-long videos. We construct a dataset of 500K samples by combining our high-quality long-form dataset (durations of 2-30+ minutes) with some samples from Stage 2. To ensure stability, we freeze the ViT and exclusively fine-tune the MLP projector and LLM backbone with a reduced learning rate of 2e-6. Additionally, we incorporate a mixture of 10% text-only instructions into the training dataset to preserve the general instruction-following capabilities of Video LLM.
D.2 Chain of Thought Prompting Details
To unlock the complex reasoning capabilities of Video LLMs, we employ a structured, multi-stage inference pipeline designed to emulate human cognitive processes inspired by (Zhou et al., 2024). This pipeline decomposes video understanding into four distinct phases: (1) Task Prompt Routing, where the question is classified into a specific reasoning domain; (2) Reasoning Analysis, where a specialized prompt generates intermediate reasoning steps (e.g., visual scans or timelines) without immediately answering the question; (3) Self-Reflection, where the model evaluates the validity of its own analysis; and (4) Final Answer Generation, where the verified analysis is synthesized into a natural response.
The specific prompt templates used for each stage are detailed in Tables 10 through 13.
Appendix E More Experimental Results
E.1 The Role of Video Dataset Duration
This section complements Section D.1.1 by analyzing how the duration distribution of training videos affects long-context video understanding. We follow the baseline training recipe in Table 3 and construct two subsets by duration: 1min-3min and 2min-30min. For each range, we uniformly sample 20K instances and train with the same hyperparameters and input configuration, so the only variable is the duration distribution.
Table 4 shows that both duration-restricted subsets underperform the baseline on LongVideoBench and LVBench, suggesting that reducing diversity by sampling from a single duration range can hurt generalization. Meanwhile, the 2min-30min subset achieves clear gains over both the baseline and the 1min-3min subset on long-video robustness, most notably on Video-MME (62.1 to 64.2) and also improves MVBench (58.2 to 58.5). In contrast, the 1min-3min subset consistently degrades across long-video and perception benchmarks.
Long-duration data (2min-30min) is beneficial for long-context robustness (especially Video-MME), likely because it exposes the model to longer-range temporal dependencies. However, using only long-duration videos can slightly reduce performance on some long-video QA benchmarks (e.g., LongVideoBench 62.0 to 61.6), indicating that duration-only filtering may trade off data diversity and task coverage. Overall, the results support using long-duration data as an important component, while keeping mixed-duration data.
Model #F LongVideoBench LVBench MVBench Video-MME
Baseline 64 62.0 43.4 58.2 62.1
1min-3min 64 59.8 43.1 57.4 62.2
2min-30min 64 61.6 43.5 58.5 64.2
Table 4: Comparison with Different Video Dataset Duration.
E.2 Multi-Stage Training
This section empirically validates the efficacy of the multi-stage curriculum detailed in Section D.1.2 by evaluating checkpoints at the end of each training stage. Table 5 presents the quantitative progression of the model performance.
[…]
Stage #F Video-MMMU LongVideoBench LVBench EgoSchema MVBench $_{\text{dev}} MLVU$$_{\text{test}}$ Video-MME Avg
Baseline 32 43.3 58.6 36.4 57.9 53.1 59.4 44.1 56.4 51.2
1 32 46.0 59.5 42.6 60.5 59.2 64.4 49.5 61.8 55.4
2 64 44.9 62.0 43.4 63.7 58.2 68.2 53.4 62.1 56.9
3 64 46.6 61.4 44.7 64.5 59.4 69.2 53.4 64.2 57.9
Table 5: Stage Performance Comparison.
E.3 CoT Inference Visualizations
We present qualitative examples of VidLaDA’s reasoning under the Chain-of-Thought framework, accelerated by MARS-Cache. Figure 8 displays the task, where the model utilizes intermediate temporal analysis to correctly reconstruct the chronological order of disjoint scenes. Figure 9 illustrates the scenario, where the model successfully deduces the subject’s expertise level by linking specific visual adjustments to causal intent.
E.4 Ablation Studies Details
In this section, we provide a comprehensive analysis of the proposed MARS-Cache. We validate our design choices regarding anchor token searching, modality-wise, and layer-wise refresh rates. Unless otherwise stated, all ablation experiments are conducted on EgoSchema (subset) and MLVU (test) to evaluate reasoning and long-context capabilities, respectively. The model layers are simply divided into four groups: Group 1 (layers 0-7), Group 2 (8-15), Group 3 (16-23), and Group 4 (24-31). The detailed settings for the ablation study are provided in Appendix F.5.
E.4.1 Effectiveness of Anchor Token Searching
We first investigate the necessity of preserving global connectivity via anchor tokens and identify which network depths benefit most from this mechanism. We compare our anchor token searching against the baseline with full attention (“-”) and variants where anchor tokens are removed (i.e., “0”, frame-wise chunk attention).
Table 6 validates the necessity of anchor tokens, as removing them completely (Row 2) significantly degrades performance compared to the baseline. Regarding token allocation, we find that a tapered strategy assigning more tokens to shallow layers and fewer to deep layers yields optimal results (Last Row). Furthermore, maintaining full attention in shallow layers to preserve structural features while applying aggressive anchor pruning in deep layers (Row 4) achieves competitive performance with similar FLOPs. This suggests that deep semantic layers can be heavily compressed provided that early structural information remains intact.
Table 6: Ablation on Count of Anchor Token. We compare the retention of anchor tokens across different layer groups. “-” indicates full attention (baseline), “0” indicates no anchor tokens (pure chunk attention), and numbers indicate the count of retained anchor tokens. FLOPs refers to the FLOPs of attention calculation only.
G1 G2 G3 G4 Ego MLVU FLOPs
- - - - 65.0 42.6 100%
0 0 0 0 63.8 40.7 53%
32 32 32 32 65.0 41.5 66%
- - 32 32 65.2 45.3 83%
128 128 128 128 64.6 44.4 93%
128 32 32 32 65.6 41.8 73%
- 32 32 32 65.8 41.7 75%
- 128 64 32 66.8 44.0 84%
E.4.2 Modality-Wise Refreshing Strategy
A main premise of MARS-Cache is that visual representations are temporally more stable than textual hidden states during the decoding process. Consequently, the visual cache can be refreshed less frequently. We define the Vision/Text Refresh Ratio ($R_{v/t}$) as the ratio of their respective update intervals.
Figure 7 presents the results under baseline text refresh intervals ($\tau_{(*,t)}=16$). We observe a consistent trend: increasing the refresh interval for vision tokens (i.e., $R_{v/t}>1$) maintains or even slightly degrades accuracy while boosting Throughput (TPS). Specifically, with $\tau_{(*,t)}=16$, the ratio of around 2 achieves the optimal balance. Pushing the ratio too high (e.g., $R_{v/t}>2$) eventually degrades performance, indicating that while visual states drift slowly, they are not static.
Figure 7: Impact of Vision/Text Refresh Ratios. We test varying the visual cache refresh interval while keeping the text interval fixed at 16.
E.4.3 Layer-wise Refreshing Strategy
Based on the observation that deep layers exhibit higher hidden state drift than shallow layers, we can utilize a hierarchical schedule where update frequencies increase with network depth.
Table 7 presents the results under this strategy. The results demonstrate that the pyramid refresh schedule (decreasing intervals from Group 1 to Group 4) yields the best performance. The configuration $64\rightarrow 32\rightarrow 16\rightarrow 8$ (Row 6) achieves a high score of 68.0 on EgoSchema, and the second-best score 45.3 on MLVU, outperforming uniform schedules. Furthermore, when combining this with the Modality-Wise strategy (where $R_{v/t}=2$), we see further robustness. This confirms that the two refresh strategies are orthogonally efficient.
Table 7: Ablation on Layer-wise Refreshing. We compare various hierarchical schedules. The cell background color indicates the update frequency: light green denotes low-frequency updates (e.g., interval 64), while dark green denotes high-frequency updates (e.g., interval 4/8).
G1 G2 G3 G4 Ego TPS MLVU TPS
Uniform Modality Refresh ($\tau_{(g,v)}=\tau_{(g,t)}$)
32 32 32 32 65.0 9.1 42.6 9.3
64 64 32 32 64.8 10.3 41.2 10.3
32 32 16 16 66.6 7.8 44.5 7.8
32 32 16 8 67.0 6.7 45.9 6.8
32 16 8 4 66.6 4.6 44.8 4.7
64 32 16 8 68.0 7.0 45.3 7.0
Modality-Aware Refresh ($\tau_{(g,v)}=2\tau_{(g,t)}$)
32 32 16 8 67.0 6.7 45.9 6.8
($\times 2$) 66.8 9.2 46.9 9.3
32 16 8 4 66.6 4.6 44.8 4.7
($\times 2$) 66.4 6.8 44.3 7.0
64 32 16 8 68.0 7.0 45.3 7.0
($\times 2$) 65.8 9.4 45.3 9.5
E.4.4 Overhead of Anchor Token Searching
We analyze the overhead introduced by the Adaptive Anchor Token Searching mechanism. Specifically, we evaluate the impact of the number of sampled query tokens used to compute the proxy attention matrix. As shown in Table 8, using a very small number of tokens (16) for vision and text leads to unstable anchor token identification and lower performance. Conversely, using the full sequence to compute the attention score map incurs heavy computational and memory costs (OOM). We find that setting the query set size to 128 provides a robust trade-off between identifying high-quality global sink nodes and maintaining low search overhead.
Table 8: Analysis of Search Token Overhead. We measure the impact of the number of query tokens used for proxy attention on model performance and computational cost. GFLOPs and Memory indicate the overhead for a single calculation.
#Vis #Txt Ego MLVU GFLOPs Memory(MB)
16 16 65.0 42.6 7.1 82.6
32 32 64.8 43.9 14.1 165.2
64 64 64.2 44.0 28.2 330.3
128 128 65.2 45.1 56.4 660.6
256 256 66.6 44.0 112.7 1351.7
$|V|$ $|T|$ - - - OOM
E.5 Causal Inference Visualizations
To explore the potential of DLM-based Video LLMs in modeling complex event dependencies, we conduct a preliminary visualization case study following the methodology of (Chen et al., 2024b). As illustrated in Figure 10, the DLM-Based Video LLM, VidLaDA, demonstrates the capability to identify non-local causal links (green arrows) bridging the temporal gap, whereas the AR baseline (LLaVA-OneVision) misses these dependencies. This demonstrates that DLM-based Video LLMs, leveraging their inherent bidirectional modeling capabilities, may possess significant potential to achieve robust performance in complex causal inference tasks.
Appendix F Experimental Details
F.1 Experimental Setup
Implementation Details. VidLaDA is implemented using the LLaDA-8B (Nie et al., 2025) as the LLM backbone, coupled with SigLIP2-SO400M (Tschannen et al., 2025) as the ViT, following the LLaDA-V (You et al., 2025). The model is trained following the three-stage curriculum strategy described in Section D.1. The specific training settings are listed in Table 9. Training is conducted on 32 NVIDIA H200 GPUs with a global batch size of 64. The learning rate can refer to Table 3.
Table 9: Training Setting Details for VidLaDA.
Item Value
Total Batch Size 64
Warmup Ratio 3%
Optimizer AdamW
LR Scheduler Cosine
Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.999
Adam $\epsilon$ 1e-8
Weight Decay 0
Gradient clipping 1.0
Datasets and Evaluation Benchmarks. We conduct a comprehensive evaluation across eight diverse benchmarks to assess VidLaDA’s capabilities in video understanding. To rigorously test long-term temporal dependencies and understanding, we employ Video-MME (Fu et al., 2025) (w/o subtitles), MLVU (Zhou et al., 2025) (reporting both Dev and Test splits), LVBench (Wang et al., 2025b), and LongVideoBench (Wu et al., 2024). These datasets challenge models with extended video contexts ranging from minutes to hours, requiring sustained attention over long sequences. For fine-grained temporal perception, we utilize MVBench (Li et al., 2024b), which comprises 20 distinct tasks requiring precise action and attribute recognition. Furthermore, we evaluate high-level reasoning and expert domain knowledge using EgoSchema (Mangalam et al., 2023), a benchmark for complex egocentric video understanding, and Video-MMMU (Hu et al., 2025), which tests multi-disciplinary professional understanding.
F.2 Intra-Frame Understanding on MME
To empirically validate the spatial robustness of VidLaDA compared to Autoregressive baselines, as discussed in Section 3.1.2, we conduct a controlled spatial permutation experiment on the MME benchmark (Yin et al., 2024). For quantitative evaluation, we report the MME Sum score, which aggregates performance across the Perception and Cognition subtasks.
We employ an information-guided shuffling protocol to test the sensitivity of the model to the position of semantic features. First, we compute the $\ell_{2}$-norm of the output features from the vision encoder for every patch, serving as a proxy for information density. We then identify the top-$k$ tokens with the highest norms, setting $k=256$, to isolate the most semantically salient regions of the image. To manipulate the sequence order, we define a position ratio $r\in[0,1]$. The input sequence is constructed such that the selected high-norm tokens are placed starting at index $\lfloor r\times(|V|-k)\rfloor$, where $|V|$ is the total number of visual tokens. Consequently, $r=0$ positions the high-information tokens at the visual start, while $r=1$ pushes them to the visual end, with the remaining low-norm tokens filling the rest of the sequence. By evaluating the performance variance across different $r$ values, we can determine whether a model exhibits position-dependent early receptive field biases.
F.3 Inter-Frame Interaction on RexTime
We evaluate variable temporal position reasoning using the ReXTime benchmark (Chen et al., 2024a), utilizing videos sourced from the QVHighlights (Lei et al., 2021) dataset. We conduct two distinct experiments to analyze the impact of event temporal position and sparsity in Section 3.1.2.
To analyze how the position of a critical event within a video affects model performance (corresponding to Figure 2(b)), we categorize the test samples based on the normalized temporal timestamp of the event. Let $T_{V}$ be the total duration of the video and $t_{\text{start}},t_{\text{end}}]$ be the ground-truth time interval of the relevant event. We define the Event Position Ratio as $r=(t_{\text{start}}+t_{\text{end}})/(2\cdot T_{V})$. We partition the test set into 15 uniform bins. We evaluate the model’s accuracy within each bin using a uniform frame sampling strategy with $\{8,32,64\}$ frames. This setup reveals potential biases in processing information located at different temporal stages (start, middle, or end) of the sequence.
For the sparsity analysis (corresponding to Figure 2(c)), we examine the model’s robustness to sparse visual evidence by progressively reducing the number of sampled key frames ($N_{\text{key}}$). Specifically, we vary $N_{\text{key}}$ in decreasing order (from 16 down to 1) to observe performance stability under constrained evidence. Furthermore, we introduce a variable number of non-key context frames ($N_{\text{noise}}$), for every fixed $N_{\text{key}}$, we iterate through a range of noise levels where the number of non-key frames is sampled from the set $\{0,2,4,\dots,32\}$. The final accuracy reported for a specific $N_{\text{key}}$ is the mean accuracy calculated across all these $N_{\text{noise}}$ settings, with the standard deviation illustrated as the shaded region in the figure.
F.4 CoT Inference Settings
This section details the inference hyperparameters employed for the reasoning experiments using the Chain-of-Thought (CoT) framework described in Appendix D.2. Regarding visual input processing, we uniformly sample up to 32 frames from the input video.
For the intermediate reasoning analysis, we allocate a generation budget of 1024 tokens for LongVideoBench and the EgoSchema (subset), and 2048 tokens for MLVU (test), to accommodate detailed thought processes, and we report the best performance achieved. Furthermore, capitalizing on the non-autoregressive nature of the underlying Diffusion Language Model, we enable parallel decoding (Wu et al., 2025) to accelerate the inference of these extended reasoning chains. Additionally, for the MLVU benchmark, we also employ CreditDecoding (Wang et al., 2025a) to ensure decoding stability.
Specifically, we adopt left-to-right block-wise decoding, a prevalent technique in current DLM inference that generates multiple tokens simultaneously within a sliding window (Nie et al., 2025; Ye et al., 2025c; You et al., 2025), and configure the process with a block length of 64.
F.5 Ablation Settings
Distinct from the multi-stage Chain-of-Thought framework described in Appendix F.4, our ablation studies prioritize experimental control and throughput stability to rigorously evaluate the efficiency of the proposed MARS-Cache. To decouple the efficiency metrics from the variance introduced by dynamic task routing and variable-length generation, we adopt an explicit thought-tagging strategy inspired by LLaVA-CoT (Xu et al., 2025).
Specifically, we utilize a single-turn prompt design that instructs the model to enclose its reasoning process within specific tokens. Leveraging the unique capability of Diffusion Language Models to manipulate the target sequence initialization, we explicitly inject these start and end thinking tags into the masked sequence at predefined positions during the forward process. This constraint strictly enforces a fixed thinking generation length, thereby eliminating length-induced latency variations and ensuring that the reported Tokens Per Second (TPS) metrics reflect the pure algorithmic efficiency of the decoding strategy. Under this controlled protocol, we measure efficiency in terms of FLOPs and Throughput (TPS) on a single NVIDIA H200 GPU, configured with a generation length of 128, 128 diffusion steps, and a block length of 32, without employing parallel decoding.
Figure 8: Reasoning Analysis. An inference example of VidLaDA via MARS-Cache on a case of Dynamic Action & Temporal Evolution.
Figure 9: Reasoning Analysis. An inference example of VidLaDA via MARS-Cache on a case of Complex Logic & Causal Reasoning.
(a) Ground Truth Causal Graph.
(b) Causal Inference from VidLaDA.
(c) Causal Inference from LLaVA-OV.
Figure 10: Causal Inference Examples. The blue arrows indicate the chronological video order. Dashed yellow arrows denote the ground truth causal dependencies. Green arrows represent correctly predicted links, while red arrows and crosses mark missed dependencies. VidLaDA successfully models long-range causal links, whereas the AR baseline fails to connect distant events.
### Stage 1: Task Prompt Routing
ROUTER_TEMPLATE = ”””
# Role
You are an expert Task Router for a Video Understanding AI System.
Your objective is to select the single best reasoning strategy (Category) for the User’s Question.
# Input Data
- Video Duration: {duration}
- User Question: ”{question}”
# Evaluation Criteria (The Strategies)
1. **Static Entity & Spatial Perception**:
    - Focus: Static visual details in a single frame (color, object count, OCR text, relative position).
    - Use when: The question is about _visible properties_ regardless of time.
2. **Dynamic Action & Temporal Evolution**:
    - Focus: Movements, action sequences, speed, trajectories, or changes over time.
    - Use when: The question contains dynamic verbs or asks about the _flow of events_.
3. **Complex Logic & Causal Reasoning**:
    - Focus: Inferred meaning, intentions, ”Why”, emotions, social interactions, or counterfactuals.
    - Use when: The answer is not explicitly visible and requires deep thinking.
4. **Fine-grained Retrieval & Temporal Grounding**:
    - Focus: Locating a specific timestamp, frame, or ”needle” in a haystack.
    - Use when: The question asks ”When”, ”Find the moment”, or for a specific detail in a long video.
5. **Global Summarization & Holistic Understanding**:
    - Focus: Summaries, titles, main themes, or condensing the video content.
    - Use when: The user wants a high-level overview.
6. **General / Uncertainty / Fallback (Default Mode)**:
    - Focus: Conversational queries, simple greetings, vague questions, or **when you are UNSURE**.
    - Use when:
    1. The question does not clearly fit 1, 2, 3, 4, or 5.
    2. The question is ambiguous.
    3. You feel low confidence in assigning a specialized category.
    - **Strategy**: It is better to choose 6 than to force a wrong specialized category.
# Decision Logic & Rules
1. Analyze the question’s intent against the keywords of 1-5.
2. **The ”Uncertainty Rule”**: If the question seems to straddle multiple categories significantly, or if it lacks specific keywords to trigger 1-5, **IMMEDIATELY CHOOSE 6**.
3. Do not over-interpret simple questions. For example, ”What is this?” is typically 6 (General), unless it asks ”What action is this?” (2).
# Output
Return ONLY the single integer of the category (1, 2, 3, 4, 5, or 6).
Table 10: The Prompt Template for Stage 1: Task Prompt Routing. The model classifies the user query into one of six distinct reasoning domains.
### Stage 2: Reasoning Analysis (Visual, Temporal, Logic)
PROMPT_STATIC_SPATIAL_ANALYSIS = ”””
# Role
You are a Computer Vision Specialist focusing on fine-grained visual details.
# Input
User Question: ”{question}”
# Task
Based strictly on the User Question above, perform a ”Visual Scan” to extract relevant visual evidence.
*Focus ONLY on the specific objects, attributes, or text mentioned or implied by the question.** DO NOT answer the question directly yet.
# Visual Scan Instructions
1. **Target Identification**: Identify the key objects or regions specifically requested in the question.
2. **Attribute Verification**: Analyze the color, shape, count, or text (OCR) of these specific targets.
3. **Spatial Check**: Describe the relative position of these targets if relevant to the question.
4. **Filtering**: Ignore visual details that are irrelevant to the user’s specific inquiry.
# Output Format
[Visual Scan]: <<Provide detailed observations strictly related to the question targets.>>”””
PROMPT_DYNAMIC_TEMPORAL_ANALYSIS = ”””
# Role
You are an Action Analysis Expert.
# Input
User Question: ”{question}”
# Task
Construct a ”Temporal Timeline” for the specific actions or events queried by the user.
*Filter out background activities that do not help answer the specific question.**
DO NOT answer the question directly yet.
# Timeline Instructions
1. **Scope Definition**: Identify which specific actor or event the question is asking about.
2. **Action Segmentation**: Log the start and end of only the relevant actions.
3. **Dynamics Analysis**: Describe the speed, direction, or intensity specifically for the queried action.
4. **Sequence Verification**: Ensure the chronological order matches the events mentioned in the question.
# Output Format
[Temporal Timeline]: <<Provide a step-by-step breakdown of the relevant actions.>>”””
PROMPT_COMPLEX_LOGIC_ANALYSIS = ”””
# Role
You are a Cognitive Analyst.
# Input
User Question: ”{question}”
# Task
Generate a ”Reasoning Chain” to bridge the visual gap for the specific question.
*Your goal is to find the ”Why” or ”How” behind the specific subject raised in the question.**
DO NOT answer the question directly yet.
# Reasoning Chain instructions
1. **Question Analysis**: What is the core mystery or implicit relationship in the question?
2. **Evidence Linking**: Connect specific visual cues (expressions, items) directly to the question’s intent.
3. **Causal Inference**: Deduce the cause or motivation specifically explaining the queried event.
4. **Contextual Exclusion**: Discard reasoning paths that do not help solve the specific question.
# Output Format
[Reasoning Chain]: <<Provide step-by-step logical deductions leading towards the answer.>>”””
Table 11: Reasoning Analysis Prompts (Part I). These prompts are selected if the Router outputs Category 1, 2, or 3, focusing on static perception, dynamic action, or complex logic, respectively.
### Stage 2: Reasoning Analysis (Retrieval, Summarization, Fallback)
PROMPT_RETRIEVAL_GROUNDING_ANALYSIS = ”””
# Role
You are a Precision Video Retrieval Assistant.
# Input
User Question: ”{question}”
# Task
Locate the specific moment or detail requested in the question.
*Your search must be strictly limited to the criteria defined in the question.**
DO NOT answer the question directly yet.
# Retrieval Strategy
1. **Keyword Mapping**: Map the question’s keywords to visual or auditory signatures (the ”Target”).
2. **Focused Scanning**: Scan the video specifically looking for these signatures.
3. **Timestamping**: Pinpoint the exact moments where the queried event occurs.
4. **Match Verification**: Confirm that the found segment perfectly matches the question’s description.
# Output Format
[Target Details]: <<Describe the found evidence and its specific location/time.>>”””
PROMPT_GLOBAL_SUMMARIZATION_ANALYSIS = ”””
# Role
You are a Professional Content Editor.
# Input
User Question: ”{question}”
# Task
Extract ”Synopsis Elements” necessary to satisfy the user’s specific request.
*Note: Do not just summarize the whole video. Summarize the aspect requested by the user.**
DO NOT answer the question directly yet.
# Synopsis Generation
1. **Topic Alignment**: Determine which aspect of the video (plot, theme, character arc) the user is asking about.
2. **Key Moment Extraction**: Select only the turning points that contribute to answering the question.
3. **Flow Description**: Describe the progression (Intro to Conclusion) relative to the user’s topic.
4. **Noise Reduction**: Omit plot details that are unrelated to the specific question.
# Output Format
[Global Synopsis]: <<Provide structured content elements focused on the query.>>”””
PROMPT_GENERAL_FALLBACK_ANALYSIS = ”””
# Role
You are a helpful AI Video Assistant.
# Input
User Question: ”{question}”
# Task
Provide a set of ”General Observations” that will help answer the specific question above.
*Focus your observation on the entities or concepts mentioned in the question.**
DO NOT answer the question directly yet.
# Input
Question: {question}
# Output Format
[Observations]: <<List key visual details strictly relevant to answering the question.>>”””
Table 12: Reasoning Analysis Prompts (Part II). These prompts are selected if the Router outputs Category 4, 5, or 6, handling specific retrieval, summarization, or general fallback scenarios.
### Stage 3: Self-Reflection
PROMPT_REFLECTION_BOOLEAN = ”””
# Role
You are a Quality Assurance Filter.
# Task
According to the video, evaluate if the provided **Analysis** is relevant, consistent, and helpful to answer the **User Question**.
# Input
1. User Question: ”{question}”
2. Analysis: ”{analysis_text}”
# Evaluation Rules
- Return **YES** if the analysis is specific, relevant, and logically sound.
- Return **NO** if the analysis is irrelevant, hallucinatory, contradictory, or empty.
# Output Constraint
Output ONLY the word ”YES” or ”NO”. Do not answer the question, or provide any explanation, or markdown formatting.
### Stage 4: Final Answer Generation
PROMPT_FINAL_WITH_CONTEXT = ”””
# Task
Answer the user’s question based on the provided **Video** and the **Expert Analysis** below.
# Instructions
1. **Synthesize**: Use the ”Expert Analysis” as a reliable guide to interpret the video content. It highlights the key moments or details relevant to the question.
2. **Answer**: Provide a direct, clear, and natural answer to the question.
3. **Refinement**: Do not explicitly say ”Based on the analysis…” or ”The expert mentioned…”. Just answer the question naturally as if you observed these details yourself.
# Input Data
1. **Expert Analysis (Context)**:
”{analysis_text}”
2. **User Question**: ”{question}”
# Output Format
[Answer]: <<Your final, concise answer>>
# Inference Pseudo-Code
category = model.generate(ROUTER_TEMPLATE.format(question=user_q))
expert_prompt = PROMPT_MAP[category]
analysis = model.generate(expert_prompt.format(question=user_q))
is_valid = model.generate(PROMPT_REFLECTION_BOOLEAN.format(analysis=analysis))
final_ans = Final_Gen(model, analysis, is_valid, PROMPT_FINAL_WITH_CONTEXT)
Table 13: Prompts for Self-Reflection (Stage 3) and Final Answer Generation (Stage 4), followed by the overall logic of the inference pipeline. The reflection step ensures that only high-quality intermediate reasoning is used for the final response.